# Animatable Gaussians: Learning Pose-dependent Gaussian Maps for High-fidelity Human Avatar Modeling

### 27 Nov 2023 v1

@@ 2.1 Mesh-based Human Avatars
The polygon mesh is the most popular 3D representation for its compatibility with traditional rendering pipelines. To model animatable human avatars using meshes, early approaches propose to reconstruct a character-specific textured mesh and animate it by physical simulation [65, 17] or retrieval from a database [82]. Recently, researchers tend to utilize neural networks to model dynamic textures and motions. Bagautdinov et al. [5], Xiang et al. [80, 81] and Halimi et al. [22] reconstruct topology-consistent meshes from dense multi-view videos and learn the dynamic texture in a UV space. DDC [20] and HDHumans [21] learn the deformation parameterized by both skeletons and embedded graph [69] of a pre-scanned template. DELIFFAS [36] employs DCC as a deformable template and parameterizes the light field around the body onto double surfaces for fast synthesis. These mesh-based methods require dense reconstruction, non-rigid tracking, or pre-scanned templates for representing dynamic humans. Besides, some works optimize the non-rigid deformation upon SMPL [43] from a monocular RGB [4, 3, 91] or RGB-D [7, 33] video, but the avatar quality is limited by the SMPL+D representation.
@@ 3.2 Overview
Figure 3: Illustration of the avatar representation. The driving pose is first used to generate posed position maps, which serve as the pose conditions. Then the StyleUNet takes the position maps as input and predicts pose-dependent Gaussian maps. We extract canonical 3D Gaussians from the Gaussian maps and deform them to the posed space by LBS. Finally, a high-fidelity synthesized avatar image is rendered through splatting-based rasterization.
@@ 3.3 Avatar Representation
where $\mathcal{G}_{\text{f}}(\mathbf{\Theta})$ and $\mathcal{G}_{\text{b}}(\mathbf{\Theta})$ are front and back pose-dependent Gaussian maps, respectively, and each pixel represents a 3D Gaussian [32] including a position, covariance, opacity and color. We also modulate the output color attributes on Gaussian maps with a view direction map $\mathcal{V}$ to model view-dependent variance like NeRF-based approaches [55]. We extract canonical 3D Gaussians inside the template mask from the pose-dependent Gaussian maps. As illustrated in Fig. 3, the canonical 3D Gaussians involve pose-dependent non-rigid deformations and appearances, e.g., garment wrinkles.
@@
where $\mathbf{R}$ and $\mathbf{t}$ are the rotation matrix and translation vector calculated with the skinning weights of each 3D Gaussian. It is worth mentioning that despite utilizing only front and back views for parameterizing the 3D Gaussians, the resulting point clouds still cover the side regions of the human body as demonstrated in the posed 3D Gaussians in Fig. 3. This is achievable because the positions of the pose-dependent 3D Gaussians are learnable and will be optimized to encompass the side regions. Finally, we render the posed 3D Gaussians to a desired camera view through splatting-based rasterization (Eq. 2).
@@ 4 Experiments
Dataset. We mainly utilize two public datasets for the experiments, including 3 sequences with 24 views from THuman4.0 dataset [93] and 5 sequences with 160 views from ActorsHQ dataset [25] (but we only use 47 full-body views for avatar modeling). For the comparison with AvatarReX [94], we also request their dataset including 2 sequences with 16 views. THuman4.0 and AvatarReX datasets also provide the SMPL-X [53] registrations. We fit SMPL-X for ActorsHQ dataset using an open-source tool [1]. We split each sequence as training and testing chunks, and the training chunk contains 2000 $\sim$ 3000 frames.
@@ 4.1 Comparison
Full-body Avatars. Full-body avatars including TotalSelfScan [12], X-Avatar [64] and AvatarReX [94] can realize expressive control of the body, hands and face. TotalSelfScan reconstructs full-body avatars from monocular self-rotation videos, and only displays animations that appear very rigid. X-Avatar requires 3D human scans under different poses as input for creating avatars. AvatarReX is the most relevant work with our method, i.e., creating avatars from multi-view videos. Since AvatarReX does not release the data and code, we request the dataset and results from the authors for comparison. Fig. 6 shows the comparison with AvatarReX on both training and novel poses. Fig. 6 (a) demonstrates that our method can reconstruct more faithful and vivid details compared with AvatarReX. Although AvatarReX introduces local feature patches to encode more details, it remains constrained by the representation ability of the conditional NeRF MLPs. Fig. 6 (b) shows that given a novel pose, our method not only generates more realistic details but also produces more reasonable non-rigid deformation, particularly for long dresses, in comparison with AvatarReX. This is attributed to the ability of our method to learn pose-dependent deformations on a character-specific template that has already modeled the basic shape of the wearing garments. In contrast, AvatarReX learns sparse node translations on the naked SMPL model, resulting in artifacts for long dresses. Tab. 2 reports the quantitative comparison on training pose reconstruction as ground-truth images are only available for training poses. Our method also outperforms AvatarReX on the reconstruction accuracy.
@@
Figure 8: Comparison between representations with 2D CNNs and MLPs on training pose reconstruction. (a,d) Results of the MLP baseline. (b,e) Our results. (c,f) Ground truth.
@@ 4.2 Ablation Study
2D CNNs vs. MLPs. To demonstrate the superior representation ability of 2D CNNs (StyleUNet in our settings), we replace the 2D networks with a coordinate-based MLP. The MLP takes a canonical point and pose vector as input, and returns the 3D Gaussian attributes of this point. Fig. 8 shows the animation results of our method with 2D CNNs and the baseline with MLPs. It demonstrates that 2D CNNs are able to regress more detailed and realistic appearances, while MLPs suffer from limited representation ability, yielding blurry animation results.
@@ Appendix A Implementation Details
Network Architecture. The network module in our avatar representation is a StyleUNet [73], a conditional StyleGAN-based [29] generator. Differently, we adapt the original StylelUNet by incorporating two decoders to predict both front & back Gaussian maps. The resolution of the input position map is $512\times 512$, and the resolution of the output Gaussian maps is $1024\times 1024$.
Training. We adopt the Adam optimizer [34] for training the StyleUNet with a learning rate of $5\times 10^{-4}$. The loss weights are set as: $\lambda_{\text{perceptual}}=0.01$, $\lambda_{\text{reg}}=0.005$. The batch size is 1, the total iteration number is 500k, and the training procedure takes about two days on one RTX 4090.
@@ Appendix B Experiment Details
Metric Evaluation. We utilize PSNR, SSIM [77], LPIPS [90] and FID [23] as the metrics for quantitative evaluations. PSNR and SSIM are computed on the entire image at the original resolution, while LPIPS and FID are computed on the largest square that covers the human body.
Comparison with Body-only Avatars. The quantitative comparison (Tab. 1 in the main paper) is conducted on the “subject00” sequence of THuman4.0 dataset [93]. The first 2000 frames are utilized as the training dataset, and the numerical results of Tab. 1 are evaluated on the rest 500 frames and the 18th camera view.
Comparison with Full-body Avatars. The quantitative comparison (Tab. 2 in the main paper) is conducted on the “white long sleeves” sequence of AvatarReX dataset [94]. Since the provided results of AvatarReX are trained on the whole sequence, we have to conduct a quantitative comparison on the reconstruction accuracy. The numerical results of Tab. 2 are evaluated on the first 500 frames and the 1st camera view.
@@ Appendix D Additional Experiments
Table B: Quantitative comparison between representations with 2D CNNs and MLPs.
@@
2D CNNs 29.3127 0.9664 0.0378 27.3143
MLPs 26.8961 0.9497 0.0650 87.0793
Quantitative Ablation Study on Parametric Template. We conduct a qualitative ablation study on the parametric template in Fig. 7 of the main paper. We also quantitatively compare the reconstructed parametric template with the naked SMPL-X model [53] on the animation accuracy in Tab. A. The numerical results are computed on the first 500 frames and the first camera view in the “long dress” sequence from AvatarReX dataset [94].
Quantitative Ablation Study on 2D CNNs. We additionally report quantitative results of 2D CNNs and MLPs in Tab. B to further prove the effectiveness of the introduction of 2D CNNs and 2D parameterization. The numerical results are computed on the first 500 frames and the 126th camera view in the “actor02_sequence1” from ActorsHQ dataset [25].
@@
Figure C: Failure cases. (a,c) Animation results by our method, (b,d) ground-truth images. Our method fails to model the motion of hairs.
@@ Appendix E Failure Cases
Our method cannot model the physical motion of components that are not driven by the body joints, e.g., the hairs, as illustrated in Fig. C since we model the whole body including clothes, hands and hairs as an entangled Gaussian representation. We leave for future work a disentangled and compositional representation for representing the dynamics of different components of the character.

### 15 Mar 2024 v2

@@ 2.1 Mesh-based Human Avatars
The polygon mesh is the most popular 3D representation for its compatibility with traditional rendering pipelines. To model animatable human avatars using meshes, early approaches propose to reconstruct a character-specific textured mesh and animate it by physical simulation [65, 16] or retrieval from a database [82]. Recently, researchers tend to utilize neural networks to model dynamic textures and motions. Bagautdinov et al. [4], Xiang et al. [80, 81] and Halimi et al. [21] reconstruct topology-consistent meshes from dense multi-view videos and learn the dynamic texture in a UV space. DDC [19] and HDHumans [20] learn the deformation parameterized by both skeletons and embedded graph [69] of a pre-scanned template. DELIFFAS [35] employs DDC as a deformable template and parameterizes the light field around the body onto double surfaces for fast synthesis. These mesh-based methods require dense reconstruction, non-rigid tracking, or pre-scanned templates for representing dynamic humans. Besides, some works optimize the non-rigid deformation upon SMPL [42] from a monocular RGB [3, 2, 92] or RGB-D [6, 32] video, but the avatar quality is limited by the SMPL+D representation.
@@ 3.2 Overview
Figure 3: Illustration of the posed position maps.
@@ 3.3 Avatar Representation
where $\mathcal{G}_{\text{f}}(\mathbf{\Theta})$ and $\mathcal{G}_{\text{b}}(\mathbf{\Theta})$ are front and back pose-dependent Gaussian maps, respectively, and each pixel represents a 3D Gaussian [31] including a position, covariance, opacity and color. We also modulate the output color attributes on Gaussian maps with a view direction map $\mathcal{V}$ to model view-dependent variance like NeRF-based approaches [54]. We extract canonical 3D Gaussians inside the template mask from the pose-dependent Gaussian maps. It is worth mentioning that despite utilizing only front and back views for parameterizing the 3D Gaussians, the resulting point clouds still cover the side regions and hands of the human body as demonstrated in Fig. 4. The reason is that the projection to front & back views is orthographic, thus there exist sufficient 3D Gaussians to model these parts.
Figure 4: Canonical 3D Gaussians on side regions and hands.
@@
where $\mathbf{R}$ and $\mathbf{t}$ are the rotation matrix and translation vector calculated with the skinning weights of each 3D Gaussian. Finally, we render the posed 3D Gaussians to a desired camera view through splatting-based rasterization (Eq. 2).
@@ 4 Experiments
Dataset. We mainly utilize two public datasets for the experiments, including 3 sequences with 24 views from THuman4.0 dataset [94] and 5 sequences with 160 views from ActorsHQ dataset [24] (but we only use 47 full-body views for avatar modeling). For the comparison with AvatarReX [95], we also request their dataset including 2 sequences with 16 views. THuman4.0 and AvatarReX datasets also provide the SMPL-X [52] registrations. We fit SMPL-X for ActorsHQ dataset using [91]. We split each sequence as training and testing chunks, and the training chunk contains 1500 $\sim$ 3000 frames.
@@ 4.1 Comparison
Full-body Avatars. Full-body avatars including TotalSelfScan [11], X-Avatar [64] and AvatarReX [95] can realize expressive control of the body, hands and face. TotalSelfScan reconstructs full-body avatars from monocular self-rotation videos, and only displays animations that appear very rigid. X-Avatar requires 3D human scans under different poses as input for creating avatars. AvatarReX is the most relevant work with our method, i.e., creating avatars from multi-view videos. Fig. 7 shows the comparison with AvatarReX on both training and novel poses. Fig. 7 (a) demonstrates that our method can reconstruct more faithful and vivid details compared with AvatarReX. Although AvatarReX introduces local feature patches to encode more details, it remains constrained by the representation ability of the conditional NeRF MLPs. Fig. 7 (b) shows that given a novel pose, our method not only generates more realistic details but also produces more reasonable non-rigid deformation, particularly for long dresses, in comparison with AvatarReX. This is attributed to the ability of our method to learn pose-dependent deformations on a character-specific template that has already modeled the basic shape of the wearing garments. In contrast, AvatarReX learns sparse node translations on the naked SMPL model, resulting in artifacts for long dresses. Tab. 2 reports the quantitative comparison on training pose reconstruction. Our method also outperforms AvatarReX on the reconstruction accuracy.
@@
Figure 9: Comparison between representations with different backbones on training pose reconstruction.
@@ 4.2 Ablation Study
Backbones. To demonstrate the superior representation ability of 2D CNNs (StyleUNet in our settings), we replace StyleUNet with a coordinate-based MLP and a standard U-Net [61], respectively. The MLP takes a canonical point and pose vector as input, and returns the 3D Gaussian attributes of this point. While the standard U-Net replaces the StyleUNet as the backbone. Fig. 9 shows the animation results of our method with StyleUNet and the baselines with MLPs and U-Net, respectively. First, it demonstrates that 2D CNNs are able to regress more detailed and realistic appearances, while MLPs suffer from limited representation ability, yielding blurry animation results. Second, StyleUNet outperforms the standard U-Net because of the additional modules (including style modulation and “To/From-RGB” modules) inherited from StyleGAN [28]. Overall, the 2D parameterization and StyleUNet enable our method to model high-quality human dynamic appearances.
@@ 5 Discussion
Acknowledgement. This paper is supported by National Key R&D Program of China (2022YFF0902200), the NSFC project No.62125107 and the Beijing Municipal Science & Technology Z231100005923030.
@@ Appendix A Implementation Details
Network Architecture. The network in our avatar representation is composed of StyleUNet [73], a conditional StyleGAN-based [28] generator. Differently, we adapt the original StyleUNet by incorporating two decoders to predict both front & back Gaussian maps. The resolution of the input position map is $512\times 512$, and the resolution of the output Gaussian maps is $1024\times 1024$. Specifically, we utilize three different StyleUNets to output color (3-channel), position (3-channel) and other Gaussian attributes (8-channel). In the color StyleUNet, we modulate the color output with a view direction map to model view-dependent effects. Each pixel on the view direction map indicates the angle between the view direction and the template normal. The view direction map is encoded through a tiny CNN, then the encoded feature map is injected into an intermediate decoder layer of the color StyleUNet.
Training. We adopt the Adam optimizer [33] for training the StyleUNet with a learning rate of $5\times 10^{-4}$. The loss weights are set as: $\lambda_{\text{perceptual}}=0.1$, $\lambda_{\text{reg}}=0.005$. The batch size is 1, the total iteration number is 500k, and the training procedure takes about two days on one RTX 4090.
@@ Appendix B Experiment Details
Metric Evaluation. We utilize PSNR, SSIM [77], LPIPS [90] and FID [22] as the metrics for quantitative evaluations. PSNR and SSIM are computed on the entire image at the original resolution, while LPIPS and FID are computed on the cropped minimal square that covers the human body.
Comparison with Body-only Avatars. The quantitative comparison (Tab. 1 in the main paper) is conducted on the “subject00” sequence of THuman4.0 dataset [94]. The first 2000 frames are utilized as the training dataset, and the numerical results of Tab. 1 are evaluated on the rest 500 frames and the “cam18” camera view.
Comparison with Full-body Avatars. The quantitative comparison (Tab. 2 in the main paper) is conducted on the “avatarrex_zzr” sequence of AvatarReX dataset [95]. The numerical results of Tab. 2 are evaluated on the first 500 frames and the “22010710” camera view.
@@ Appendix D Additional Experiments
Table B: Quantitative comparison between representations with different backbones.
@@
StyleUNet [73] 29.3127 0.9664 0.0378 27.3143
U-Net [61] 26.4255 0.9435 0.0507 31.3838
MLP 26.8961 0.9497 0.0650 87.0793
Table C: Quantitative ablation study on pose projection.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
w Pose Proj. 24.9932 0.9285 0.0685 45.6266
w/o Pose Proj. 23.5594 0.9189 0.0792 59.9083
Quantitative Ablation Study on Parametric Template. We conduct a qualitative ablation study on the parametric template in Fig. 7 of the main paper. We also quantitatively compare the reconstructed parametric template with the naked SMPL-X model [52] on the animation accuracy in Tab. A. The numerical results are computed on the first 500 frames and the “22010708” camera view in the “avatarrex_lbn1” sequence from AvatarReX dataset [95].
Quantitative Ablation Study on Backbones. We additionally report quantitative results of StyleUNet [73], UNet [61] and MLPs in Tab. B to further prove the effectiveness of the introduction of 2D CNNs and 2D parameterization. The numerical results are computed on the 60-560 frames and the “Cam127” camera view in the “Actor02/Sequence1” from ActorsHQ dataset [24].
Quantitative Ablation Study on Pose Projection. We quantitatively ablate pose projection in Tab. C. It shows pose projection realizes more accurate animation under novel poses. The numerical results are computed on the 1109-1119, 1447, 1515 and 1578 frames where the testing poses are out of the distribution of training poses.
@@
View Number. We quantitatively and qualitatively show the animation results trained with 3 views, 6 views and 14 views in Tab. D and Fig. C. They demonstrate that our method also supports sparse-view input and can realize comparable high-fidelity results. The numerical results are evaluated on the first 500 frames and the “22070932” camera view of the “avatarrex_zzr” sequence from AvatarReX dataset [95].
Figure C: Animation results trained with different numbers of views.
Table D: Quantitative evaluation on different view numbers.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
3 Views 30.6123 0.9807 0.0306 11.3066
6 Views 30.3565 0.9803 0.0310 10.9966
14 Views 30.7622 0.9816 0.0297 10.6744
Figure D: Failure cases. (a,c) Animation results by our method, (b,d) ground-truth images. Our method fails to model the motion of hairs.
@@ Appendix E Failure Cases
Our method cannot model the physical motion of components that are not driven by the body joints, e.g., the hairs, as illustrated in Fig. D, since we model the whole body including clothes, hands and hairs as an entangled Gaussian representation. We leave for future work a disentangled and compositional representation for modeling the dynamics of different components of the character.